# Dynamical black hole entropy beyond general relativity from the Einstein frame

### v2

@@ 4.3 The comparison first law
Since the asymptotic forms of $g_{ab}$ and $g^{E}_{ab}$ agree up to a normalization factor, the mass and angular momenta of the two spacetimes agree up to a normalization factor according to the Iyer-Wald definitionsJacobson:1995uq ; Iyer:1994ys . And the angular velocities are in agreement as $\xi_{E}^{a}\propto\xi^{a}$. One can determine the relationship between $M^{E}$ and $M$ as $M^{E}=\alpha M$ since the time translations at infinity are related as $(\partial_{t^{E}})^{a}=\alpha(\partial_{t})^{a}$, and similarly we have $J^{E}=\alpha J$. Notice that $\kappa^{E}=\alpha\kappa$, $\delta S_{\text{dyn}}$ does not change to first order as we switch from $v^{E}-$identification to $v-$identification, and $S^{E}_{\text{dyn}}$ remains unchanged to first order if we use $v$ instead of $v^{E}$ in the formula. The comparison first law for an arbitrary cross-section is translated into the claim in the original frame as
@@
4.4 The second law for vacuum perturbations
For vacuum perturbations we have no external matter fields and $\delta T_{ab}=0$, so the dynamical black hole entropy is invariant to first order. In order to study the non-trivial leading-order behavior of $S_{\text{dyn}}$, we have to pay attention to the second order in perturbation theory. The first-order variation of the Raychaudhuri equation in the Einstein frame reads
@@ 4.4 The second law for vacuum perturbations
Because $\theta^{E}$ vanishes asymptotically at late times, $\delta\theta^{E}=0$ holds on $\mathcal{H}^{+}$. Vary the Raychaudhuri equation combined with the Einstein equation twice to obtain [^4] Notice that the stress-energy tensor of the auxiliary scalar field contributes to the second-order variation of $\mathcal{T}_{ab}$.
@@
First, (56) is invariant to second order in the perturbation theory as we decide to compare spacetime points sharing the same $v$ rather than $v^{E}$ on the horizon. Since by the argument similar to that of (48), switching from $v^{E}-$identification to $v-$identification for $\delta S_{\text{dyn}}$ and $\delta^{2}S_{\text{dyn}}$ leads to third-order corrections as $\partial_{v}S_{\text{dyn}}\sim\mathcal{O}(\epsilon^{2})$, and altering the identification of spacetime points also leads to third-order corrections on the right hand side of (56). Moreover, for vacuum perturbations $S^{E}_{\text{dyn}}$ is invariant to second order if we replace $v^{E}$ with $v$ in the formula. As $S^{E}_{\text{dyn}}$ is a constant to first order under the vacuum perturbations, in the perturbed spacetime
@@
where $S^{E}_{\text{BH}}=\frac{1}{4G_{D}}\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}}$ is the Bekenstein-Hawking entropy in the Einstein frame and $C$ is a constant. Requiring that $S^{E}_{\text{BH}}$ remains finite at future infinity $v^{E}\to\infty$ in the perturbed spacetime leads to the conclusion that
@@
$\begin{split}\delta\partial_{v}\ln\left[\sqrt{\gamma}f^{\prime}(R)\right]=\frac{f^{\prime}(R)\partial_{v}\delta\sqrt{\gamma}+\sqrt{\gamma}f^{\prime\prime}(R)\partial_{v}\delta R}{\sqrt{\gamma}f^{\prime}(R)}=\delta\theta+\frac{f^{\prime\prime}(R)}{f^{\prime}(R)}\partial_{v}\delta R\overset{\mathcal{H}^{+}}{=}0.\end{split}$ (59)
@@
$\begin{split}\delta\sigma^{E}_{ab}&=\frac{1}{2}\mathcal{L}_{k^{E}}\delta\gamma^{E}_{ab}-\frac{1}{D-2}\delta\theta^{E}\gamma_{ab}\\ &=\frac{1}{2}\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\mathcal{L}_{k}\delta\gamma_{ab}+\frac{1}{D-2}\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\frac{f^{\prime\prime}(\varphi)}{f^{\prime}(\varphi)}\gamma_{ab}\mathcal{L}_{k}\delta\varphi\\ &=\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}-\frac{1}{D-2}\delta\theta\gamma_{ab}\right]\\ &=\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\delta\sigma_{ab},\end{split}$ (60)
@@
$\delta\sigma^{E}_{ab}\delta\sigma_{E}^{ab}=\gamma_{E}^{ac}\gamma_{E}^{bd}\delta\sigma^{E}_{ab}\delta\sigma^{E}_{cd}=\gamma_{E}^{ac}\gamma_{E}^{bd}\left[f^{\prime}(\varphi)\right]^{\frac{4}{D-2}}\delta\sigma_{ab}\delta\sigma_{cd}=\gamma^{ac}\gamma^{bd}\delta\sigma_{ab}\delta\sigma_{cd}=\delta\sigma_{ab}\delta\sigma^{ab}.$ (61)
@@
4.5 Relation to the Wald entropy of the generalized apparent horizon
If the derivative of the Wall entropy with respect to the affine parameter $\mathrm{d}S_{\text{Wall}}/\mathrm{d}v$ is positive, which follows from the second law of black hole thermodynamicsJacobson:1995uq ; Wall:2015raa , it suggests that dynamical black hole entropy is associated to the entropy of a surface inside the black hole. Although the location of the apparent horizon associated with the prescribed cross-section $\mathcal{C}(v)$ is ambiguous as it depends on the choice of the simultaneous surface, for perturbations of a stationary black hole with bifurcate Killing horizon, the notion of the area of an apparent horizon corresponding to cross-section $\mathcal{C}(v)$ is well defined to first orderHollands:2024vbe . And it was shown in the appendix A of Hollands:2024vbe that the dynamical black hole entropy is equal to the area entropy of the apparent horizon to first order in GR. Visser and Yan also provided a more pedagogical proof of this claim using Gaussian null coordinates (GNC) systemVisser:2024pwz . Their proof can be applied to $f(R)$ gravity in the Einstein frame since it’s only a geometric fact. And in the Einstein frame this relation reads
@@ 4.5 Relation to the Wald entropy of the generalized apparent horizon
$\theta_{\lambda^{E}}=\frac{\partial}{\partial{\lambda^{E}}}\ln\sqrt{\gamma^{E}}=0,$ (66)
@@
in the original frame, where $c$ is a constant and $\lambda$ is the affine parameter of the outgoing null normal in the original frame. The generalized expansion for $f(R)$ gravity in the original frame is defined asMatsuda:2020yvl
@@
5 Dynamical black hole entropy in canonical scalar-tensor theory
@@
5.1 The conformal transformation and the Einstein frame
@@
And we assume that $F(\varphi)>0$ to avoid tensor ghostsMatsuda:2020yvl . We turn to the Einstein frame and redefine the scalar field asDicke:1961gz
$g^{E}_{ab}=\left[F(\varphi)\right]^{\frac{2}{D-2}}g_{ab},\quad\phi=\frac{1}{\sqrt{16\pi G_{D}}}\int\mathrm{d}\tilde{\varphi}\ \sqrt{F(\tilde{\varphi})^{-1}+\frac{2(D-1)}{D-2}\left[\frac{F^{\prime}(\tilde{\varphi})}{F(\tilde{\varphi})}\right]^{2}}.$ (71)
@@
And for an asymptotically flat stationary black hole solution to the scalar-tensor theory, the spacetime given by (71) remains an asymptotic flat stationary black hole with the same event horizon. Without loss of generality we assume $\varphi$ tends to be a constant at spatial infinity. So the surface gravities of the two black hole spacetimes are related as $\kappa^{E}=\alpha\kappa$ with $\alpha$ representing the value of $\left[F(\varphi)\right]^{-\frac{1}{D-2}}$ at infinity.
5.2 The physical process first law
@@ 5.3 The comparison first law
Then as $\kappa^{E}=\alpha\kappa$, $M^{E}=\alpha M$ and $J^{E}=\alpha J$ with $\alpha$ representing the inverse of the conformal factor at infinity, we obtain the comparison first law in the original frame
@@
5.4 The second law for vacuum perturbations
If external matter fields are absent, for vacuum perturbations $\delta T^{E}_{ab}=0$ if follows immediately that the dynamical black hole entropy is a constant to first order. So we have to pay attention to the second-order behavior to study the non-trivial change of the dynamical black hole entropy. The first-order variation of the Raychaudhuri equation implies that $\delta\theta^{E}$ is 0 along the horizon. We consider the second-order variation of the Raychaudhuri equation
@@ 5.4 The second law for vacuum perturbations
$\begin{split}\delta\sigma^{E}_{ab}&=\frac{1}{2}\mathcal{L}_{k^{E}}\delta\gamma^{E}_{ab}\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}+\frac{1}{D-2}\gamma_{ab}\frac{F^{\prime}(\varphi)}{F(\varphi)}\mathcal{L}_{k}\delta\varphi\right]\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}-\frac{1}{D-2}\delta\theta\gamma_{ab}\right]\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\delta\sigma_{ab}(\epsilon^{2}).\end{split}$ (83)
@@
$\frac{\kappa}{2\pi}\Delta\delta^{2}S_{\text{dyn}}=\frac{1}{4\pi G_{D}}\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa vF(\varphi)\left[\delta\sigma_{ab}\delta\sigma^{ab}+\frac{D-1}{D-2}\left(\delta\theta\right)^{2}+\frac{1}{2F(\varphi)}\left(k^{a}\partial_{a}\delta\varphi\right)^{2}\right].$ (85)
@@
5.5 Relation to the Wald entropy of the generalized apparent horizon
@@
We would like to recast this relation in the original frame. Define the generalized expansion for canonical scalar-tensor theory asMatsuda:2020yvl
@@
6 Conclusion and discussion
In this paper we first generalize the pedagogical proof of the non-stationary physical process first law to non-vacuum black holes, and give a pedagogical deduction of the second-order behavior for $S_{\text{dyn}}$ under the vacuum perturbations in GR. We then derive both the non-stationary physical process first law and the non-stationary comparison first law for $f(R)$ gravity as well as canonical scalar-tensor theory by means of conformal transformations. Our formulae for dynamical black hole entropy agree with those derived by the Noether charge method. We further study the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations in those theories, and find that second law is obeyed both in the Einstein frames and in the original frames. Moreover, we determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame.
However, our methodology can be applied only to theories that can be put into the Einstein frames via conformal transformations. While there are also many theories of gravity cannot be put into the Einstein frame. One of the future works is to present a more pedagogical proof of the physical process first law and the comparison first law for theories such as Lovelock gravity. So far our arguments for dynamical black hole entropy only dwell on the classical level. It would be important to investigate at the semi-classical level whether the entropy of the Hawking radiation of a black hole will obtain a dynamical correction term after the non-stationary perturbation.
@@ Acknowledgements.
This work is partly supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grant Nos. 12075026, 12035016, 12361141825 and 12375058.

### v4

@@ IV.3 The comparison first law
Since the asymptotic forms of $g_{ab}$ and $g^{E}_{ab}$ agree up to a normalization factor, the mass and angular momenta of the two spacetimes agree up to a normalization factor according to the Iyer-Wald definitions[25, 6]. And the angular velocities are in agreement as $\xi_{E}^{a}\propto\xi^{a}$. One can determine the relationship between $M^{E}$ and $M$ as $M^{E}=\alpha M$ since the time translations at infinity are related as $(\partial_{t^{E}})^{a}=\alpha(\partial_{t})^{a}$, and similarly we have $J^{E}=\alpha J$. Notice that $\kappa^{E}=\alpha\kappa$, $\delta S_{\text{dyn}}$ does not change to first order as we switch from $v^{E}-$identification to $v-$identification, and $S^{E}_{\text{dyn}}$ remains unchanged to first order if we use $v$ instead of $v^{E}$ in the formula. The comparison first law for an arbitrary cross-section is translated into the claim in the original frame as
@@
IV.4 The second law for vacuum perturbations
For vacuum perturbations we have no external matter fields and $\delta T_{ab}=0$, so the dynamical black hole entropy is invariant to first order. In order to study the nontrivial leading-order behavior of $S_{\text{dyn}}$, we have to pay attention to the second order in perturbation theory. The first-order variation of the Raychaudhuri equation in the Einstein frame reads
@@ IV.4 The second law for vacuum perturbations
Because $\theta^{E}$ vanishes asymptotically at late times, $\delta\theta^{E}=0$ holds on $\mathcal{H}^{+}$. Notice that the stress-energy tensor of the auxiliary scalar field contributes to the second-order variation of $\mathcal{T}_{ab}$. We vary the Raychaudhuri equation combined with the Einstein equation twice to obtain
@@
First, (56) is invariant to second order in the perturbation theory as we decide to compare spacetime points sharing the same $v$ rather than $v^{E}$ on the horizon. Since by the argument similar to that of (48), switching from $v^{E}-$identification to $v-$identification for $\delta S_{\text{dyn}}$ and $\delta^{2}S_{\text{dyn}}$ leads to third-order corrections as $\partial_{v}S_{\text{dyn}}\sim\mathcal{O}(\epsilon^{2})$, and altering the identification of spacetime points also leads to third-order corrections on the right-hand side of (56). Moreover, for vacuum perturbations $S^{E}_{\text{dyn}}$ is invariant to second order if we replace $v^{E}$ with $v$ in the formula. As $S^{E}_{\text{dyn}}$ is a constant to first order under the vacuum perturbations, in the perturbed spacetime
@@
where $S^{E}_{\text{BH}}=\frac{1}{4G_{D}}\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}}$ is the Bekenstein-Hawking entropy in the Einstein frame and $C$ is a constant. The solution to this equation reads
$S^{E}_{\text{BH}}=C+\varepsilon v^{E}+\mathcal{O}(\epsilon^{2}),$ (58)
where $\varepsilon$ is a constant. Requiring that $S^{E}_{\text{BH}}$ remains finite at future infinity $v^{E}\to\infty$ in the perturbed spacetime leads to the conclusion that
@@
$\begin{split}\delta\partial_{v}\ln\left[\sqrt{\gamma}f^{\prime}(R)\right]&=\frac{f^{\prime}(R)\partial_{v}\delta\sqrt{\gamma}+\sqrt{\gamma}f^{\prime\prime}(R)\partial_{v}\delta R}{\sqrt{\gamma}f^{\prime}(R)}\\ &=\delta\theta+\frac{f^{\prime\prime}(R)}{f^{\prime}(R)}\partial_{v}\delta R\overset{\mathcal{H}^{+}}{=}0.\end{split}$ (60)
@@
$\begin{split}\delta\sigma^{E}_{ab}&=\frac{1}{2}\mathcal{L}_{k^{E}}\delta\gamma^{E}_{ab}-\frac{1}{D-2}\delta\theta^{E}\gamma_{ab}\\ &=\frac{1}{2}\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\left[\mathcal{L}_{k}\delta\gamma_{ab}+\frac{2}{D-2}\frac{f^{\prime\prime}(\varphi)}{f^{\prime}(\varphi)}\gamma_{ab}\mathcal{L}_{k}\delta\varphi\right]\\ &=\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}-\frac{1}{D-2}\delta\theta\gamma_{ab}\right]\\ &=\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}\delta\sigma_{ab},\end{split}$ (61)
@@
$\begin{split}\delta\sigma^{E}_{ab}\delta\sigma_{E}^{ab}&=\gamma_{E}^{ac}\gamma_{E}^{bd}\delta\sigma^{E}_{ab}\delta\sigma^{E}_{cd}=\gamma_{E}^{ac}\gamma_{E}^{bd}\left[f^{\prime}(\varphi)\right]^{\frac{4}{D-2}}\delta\sigma_{ab}\delta\sigma_{cd}\\ &=\gamma^{ac}\gamma^{bd}\delta\sigma_{ab}\delta\sigma_{cd}=\delta\sigma_{ab}\delta\sigma^{ab}.\end{split}$ (62)
@@
IV.5 Relation to Wald entropy of the generalized apparent horizon
If the derivative of the Wall entropy with respect to the affine parameter $\mathrm{d}S_{\text{Wall}}/\mathrm{d}v$ is positive, which follows from the second law of black hole thermodynamics[25, 8], it suggests that dynamical black hole entropy is associated to the entropy of a surface inside the black hole. Although the location of the apparent horizon associated with the prescribed cross-section $\mathcal{C}(v)$ is ambiguous as it depends on the choice of the simultaneous surface, for perturbations of a stationary black hole with bifurcate Killing horizon, the notion of the area of an apparent horizon corresponding to cross-section $\mathcal{C}(v)$ is well defined to first order[1]. And it was shown in the appendix A of [1] that the dynamical black hole entropy is equal to the area entropy of the apparent horizon to first order in GR. Visser and Yan also provided a more pedagogical proof of this claim using Gaussian null coordinates (GNC) system[9]. Their proof can be applied to $f(R)$ gravity in the Einstein frame since it’s only a geometric fact. And in the Einstein frame this relation reads
@@ IV.5 Relation to Wald entropy of the generalized apparent horizon
$\theta^{E}_{\tilde{k}}=\frac{\partial}{\partial{\lambda^{E}}}\ln\sqrt{\gamma^{E}}=0,$ (67)
@@
in the original frame, where $c$ is a constant and $\lambda$ is the affine parameter of the outgoing null normal in the original frame. The generalized expansion for $f(R)$ gravity in the original frame is defined as[26]
@@
V Dynamical black hole entropy in canonical scalar-tensor theory
@@
V.1 The conformal transformation and the Einstein frame
@@
And we assume that $F(\varphi)>0$ to avoid tensor ghosts[26]. We turn to the Einstein frame and redefine the scalar field as[27]
$\begin{split}g^{E}_{ab}&\equiv\left[F(\varphi)\right]^{\frac{2}{D-2}}g_{ab},\\ \phi&\equiv\frac{1}{\sqrt{16\pi G_{D}}}\int\mathrm{d}\tilde{\varphi}\ \sqrt{F(\tilde{\varphi})^{-1}+\frac{2(D-1)}{D-2}\left[\frac{F^{\prime}(\tilde{\varphi})}{F(\tilde{\varphi})}\right]^{2}}.\end{split}$ (72)
@@
And for an asymptotically flat stationary black hole solution to the scalar-tensor theory, the spacetime given by (72) remains an asymptotically flat stationary black hole with the same event horizon. Without loss of generality we assume $\varphi$ tends to be a constant at spatial infinity. So the surface gravities of the two black hole spacetimes are related as $\kappa^{E}=\alpha\kappa$ with $\alpha$ representing the value of $\left[F(\varphi)\right]^{-\frac{1}{D-2}}$ at infinity.
V.2 The physical process first law
@@ V.3 The comparison first law
Then as $\kappa^{E}=\alpha\kappa$, $M^{E}=\alpha M$ and $J^{E}=\alpha J$ with $\alpha$ representing the inverse of the conformal factor at spatial infinity, we obtain the comparison first law in the original frame
@@
V.4 The second law for vacuum perturbations
If external matter fields are absent, for vacuum perturbations $\delta T^{E}_{ab}=0$ it follows immediately that the dynamical black hole entropy is a constant to first order. So we have to pay attention to the second-order behavior to study the nontrivial change of the dynamical black hole entropy. The first-order variation of the Raychaudhuri equation implies that $\delta\theta^{E}$ is 0 along the horizon. We consider the second-order variation of the Raychaudhuri equation
@@ V.4 The second law for vacuum perturbations
$\begin{split}\delta\sigma^{E}_{ab}&=\frac{1}{2}\mathcal{L}_{k^{E}}\delta\gamma^{E}_{ab}\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}+\frac{1}{D-2}\gamma_{ab}\frac{F^{\prime}(\varphi)}{F(\varphi)}\mathcal{L}_{k}\delta\varphi\right]\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\left[\frac{1}{2}\mathcal{L}_{k}\delta\gamma_{ab}-\frac{1}{D-2}\delta\theta\gamma_{ab}\right]\\ &=\left[F(\varphi)\right]^{\frac{2}{D-2}}\delta\sigma_{ab}.\end{split}$ (84)
@@
$\begin{split}\frac{\kappa}{2\pi}\Delta\delta^{2}S_{\text{dyn}}&=\frac{1}{4\pi G_{D}}\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa vF(\varphi)\\ &\times\left[\delta\sigma_{ab}\delta\sigma^{ab}+\frac{D-1}{D-2}\left(\delta\theta\right)^{2}+\frac{1}{2F(\varphi)}\left(k^{a}\partial_{a}\delta\varphi\right)^{2}\right].\end{split}$ (86)
@@
V.5 Relation to Wald entropy of the generalized apparent horizon
@@
We would like to recast this relation in the original frame. Define the generalized expansion for canonical scalar-tensor theory as[26]
@@
VI Conclusion and discussion
In this paper we first generalize the pedagogical proof of the non-stationary physical process first law to non-vacuum black holes, and give a pedagogical deduction of the second-order behavior for $S_{\text{dyn}}$ under the vacuum perturbations in GR. We then derive both the non-stationary physical process first law and the non-stationary comparison first law for $f(R)$ gravity as well as canonical scalar-tensor theory by means of conformal transformations. Our formulas for dynamical black hole entropy agree with those derived by the Noether charge method. We further study the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations in those theories, and find that second law is obeyed both in the Einstein frames and in the original frames. Moreover, we determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame.
In the future work we plan to generalize the study of dynamical black hole entropy to more general theories of gravity, such as $f(\text{Ricci})$, based on other field redefinition techniques[7, 28, 29]. So far our arguments for dynamical black hole entropy only dwell on the classical level. It would be important to investigate at the semi-classical level whether the entropy of the Hawking radiation of a black hole will obtain a dynamical correction term after the non-stationary perturbation.
@@ Acknowledgements.
This work is partly supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grants No. 12075026, No. 12035016, No. 12361141825, and No. 12375058.